# Psychologically Plausible Deep Learning with Differentiable Tversky Similarity

### v1

@@ 4 Discussion
Tversky neural networks may require task-specific hyperparameter tuning. For instance, the existence of an optimal feature bank size was suggested by our experiments. However it is unclear how to determine this number apriori. The normalization of object and prototype vectors was beneficial in TverskyResnet-50 on MNIST, but not in TverskyXORNet.
Weight tying in language models is adequate.
Our experiments with GPT-2 on PTB showed that weight tying is beneficial in both language modeling heads employing Tversky and linear projections. Our interpretation of projection layers as measurement of similarity of stimuli to prototypes explains the adequacy of weight tying in language models: Token embeddings are prototypes, and the language modeling head classifies output tokens by measuring their similarity to those prototypes. This practice with Tversky projection layers is further principled given the underlying theory of similarity.
Data-domain projection parameter specification could enable interpretable editing.
For instance, excluding handwritten sevens with the mid horizontal stroke from VisualTverskyMNISTNet’s training set would result in a prototype without that stroke, and high error rates on handwritten sevens with that stroke. The resulting model could be edited by adding a feature specifying the stroke, and editing the prototype to include the stroke. This neural network editing approach could enable the understanding and eradication of dangerous model biases.
Salience and Asymmetry:
We observed that $\alpha>\beta$ in trained Tversky neural networks (See Appendix G), implying that they weigh the distinctive features of instances more heavily than those of prototypes. We also observed that more salient stimuli exhibited more ‘goodness of form’ (Figure 3). Both of those observations are inline with Tversky’s theory, and empirical psychological observations.
Asymetry and Exploration of Semantic Relations:
Measures of object feature intersections and differences, and the saliency of objects can be computed at inference time in Tversky neural networks using Equation 3, 4, 5, and 2, which enables a more psychologically plausible exploration of semantic relations compared to the parallelogram approach (Mikolov, 2013; Pennington et al., 2014; Peterson et al., 2020), which is based on the geometric model of similarity.
@@ Appendix B Illustration of Data-Domain Visualization of Projection Layers
Appendix C Additional Artifacts from qualitative experiments
Model Params Valid ACC
MNISTNet 21 394 99.1%
VisualMNISTNet 28 394 99.1%
TverskyMNISTNet 7 023 98.7%
VisualTverskyMNISTNet 29 463 98.7%
Appendix D Training a Tversky Projection to Learn XOR with Gradient Descent
D.1 Task and Data
@@
D.2 Method
To empirically estimate the sensitivity of Tversky projection layer’s convergence to its hyperparameters, we train 12 9601296012\,96012 960 xor models consisting of the following combinations of hyperparameters:
@@ D.3 Results
Tversky projection layer successfully modeled $xor$ with as little as 1 feature. $p(conv)$ was maximal with 16 features, but not monotonously increase with feature count.
@@
Figure 8: Example of learned TverskyProjection layers optimized to model the xor function. While multiple solutions are possible (top row), some initializations do not lead to convergence (botton row). The decision boundaries were drawn as follows. Random input vectors [x,y] are uniformly sampled in the range [-2,-2] to [2,2] the tip of each vector is ploted as a colored dot, with the color representing the trained Tversky Projection layer’s output. Tversky projection layers modeling the $xor$ function, which has a boolean domain and range extend $xor$ to the real vector space. Successful models should show data points [0,0] and [1,1] in purple (class 0), and [0,1] and [1,0] in yellow (class 1)
@@
Table 4: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 6 intersection reduction methods ($A\cap B$) and 2 difference reduction methods ($A-B$). Each row corresponds to the aggregate of 3x3x2x6x9 = 972 training sessions with: 3 fbank and prototype initialization methods(orthogonal, normal, uniform), 2 normalization modes (normalized, and not normalized), 6 feature bank sizes (1, 2, 4, 8, 16, 32), and 9 random seeds. In each case, models were trained for 1000 epochs
$A\cap B$ $A-B$ n loss acc best acc p(conv)
@@
gmean ignorematch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $00.00\pm 0$ 0.00
gmean substractmatch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $00.00\pm 0$ 0.00
@@
Appendix E Tversky-GPT-2 on PTB
@@
finetuned tversky-head N N 256 163233795 nan 251.01 115.07 146.94 209.30 nan 114.71 151.94 198.93 434.14
finetuned tversky-head N N 512 163430403 nan 106.65 70.09 83.82 120.91 87.72 70.02 85.69 120.09 212.73
@@
finetuned tversky-head N N 50257 201634563 45.93 51.32 46.06 64.96 34.82 65.63 46.23 65.07 137.68 -
finetuned tversky-all-2layers N N 256 120720459 nan 454.15 247.55 410.47 426.88 - - - - -
finetuned tversky-all-1layer N N 256 113642535 nan 566.44 260.88 352.80 411.42 nan 259.32 359.88 753.23 434.32
finetuned tversky-all-2layers N N 512 120917067 191.49 313.26 198.98 223.34 314.67 - - - - -
finetuned tversky-all-1layer N N 512 113839143 188.11 287.67 186.53 202.06 296.38 nan 183.95 204.57 605.40 363.71
finetuned tversky-all-2layers N N 1024 121310283 162.13 248.66 166.77 155.31 246.82 - - - - -
@@
finetuned tversky-head Y N 256 124636419 171.24 133.77 207.72 195.19 166.38 175.55 211.08 199.08 265.12 271.04
finetuned tversky-head Y N 512 124833027 122.55 63.69 130.32 100.72 98.91 98.83 131.35 101.80 114.61 147.93
@@
finetuned tversky-all-2layers Y N 256 82123083 nan 483.32 304.70 324.18 452.82 - - - - -
finetuned tversky-all-1layer Y N 256 75045159 nan 399.90 285.73 318.80 358.06 nan 294.68 326.37 752.72 370.36
finetuned tversky-all-2layers Y N 512 82319691 193.67 311.54 222.26 232.84 325.30 - - - - -
finetuned tversky-all-1layer Y N 512 75241767 196.22 280.78 197.75 226.43 257.15 234.72 197.38 227.43 751.74 312.60
finetuned tversky-all-2layers Y N 1024 82712907 148.09 216.51 149.45 185.73 241.28 - - - - -
@@
scratch tversky-head N N 256 163233795 nan 1024.06 183.50 226.22 245.83 nan 184.45 224.82 752.90 460.48
scratch tversky-head N N 512 163430403 nan 326.21 155.25 181.24 198.25 nan 155.31 181.87 723.19 252.86
@@
scratch tversky-head N N 50257 201634563 146.50 139.44 150.07 155.97 193.20 162.25 151.26 158.19 177.91 -
scratch tversky-all-2layers N N 256 120720459 nan 802.68 178.39 278.38 269.30 - - - - -
scratch tversky-all-1layer N N 256 113642535 nan 1756.42 183.55 241.06 304.12 nan 182.73 238.18 637.42 446.48
scratch tversky-all-2layers N N 512 120917067 nan 258.01 146.35 168.63 216.07 - - - - -
scratch tversky-all-1layer N N 512 113839143 nan 294.36 143.93 171.37 234.29 nan 144.23 171.78 753.19 428.03
scratch tversky-all-2layers N N 1024 121310283 nan 181.24 129.49 135.72 174.94 - - - - -
@@
scratch tversky-head Y N 256 124636419 nan 732.28 192.21 226.59 279.30 nan 194.08 227.19 753.58 303.63
scratch tversky-head Y N 512 124833027 nan 285.05 163.16 183.53 231.27 nan 163.17 183.91 753.01 251.48
@@
scratch tversky-all-2layers Y N 256 82123083 nan 994.22 189.88 263.99 317.77 - - - - -
scratch tversky-all-1layer Y N 256 75045159 nan 2161.75 186.98 300.85 301.06 nan 187.76 282.05 752.90 657.88
scratch tversky-all-2layers Y N 512 82319691 nan 676.04 159.48 211.49 238.71 - - - - -
scratch tversky-all-1layer Y N 512 75241767 nan 290.39 156.12 184.46 239.33 nan 157.02 184.25 360.94 470.09
scratch tversky-all-2layers Y N 1024 82712907 nan 184.58 145.46 154.26 186.24 - - - - -
@@
Appendix F Tversky Resnet-50 on MNIST
Figure 10: ResNet-50 (dashed horizontal lines) and Tversky-ResNet-50 adapted (top left), fine-tuned (bottom left) and trained from scratch (bottom right) for MNIST handwritten digit classification. For Tversky-ResNet-50, feature banks of size 1 to 20 are tried. Solid horizontal lines show the best Tversky-ResNet-50 accuracy. Normalizing object, prototype and feature vectors (yellow dots) generally resulted in better accuracy than not doing so (blue dots).
@@ H.1 Resnet-50 / NABirds Experiments
H.2 Resnet-50 / MNIST Experiments
Figure 12: Wall-clock Times in seconds for [Tversky-]Resnet-50 experiments on MNIST. 4 x NVIDIA RTX A6000 were used for each run. The x-axis shows Tversky-Resnet-50 feature bank sizes. Baseline Resnet-50 are ploted at x=0.
H.3 GPT2 Experiments
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Yes, the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We do discuss limitations in the main manuscript where applicable. For instance:
    • Empirical experiments on learning Tversky Projections modeling XOR showed that some initializations may lead to convergence failure.
    • We noted that the image recognition accuracy gains observed in TverskyResnet-50 could be due to its increased parameter count. However we hypothesize that the accuracy gains are at least in part due to the use of Tversky projections, which subsequent experimental results supported.
    • Our proposed input-domain visualization method for projection parameters results in increased parameter count and induce additional training-time computation costs.
    While we agree that a separate limitation section is a good idea, in our particular case, we found that it would be more helpful to readers to address limitations inline where relevant.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We proved by construction that a single Tversky projection layer can model the xor function by providing a concrete example of such projection along its parameter values.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: This manuscript includes all the details needed to reproduce our results.
    • Our proposed parameterization of Tversky’s similarity function is fully described.
    • Our proposed Tversky Projection layer and its hyperparameters are fully described.
    • Our experimental protocols are fully specified and rely on publicly available baselines model architectures, pretrained weights, and datasets.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Our results can be reproduced from the description in the manuscript. Additionally, we will submit supplementary materials, including Pytorch implementation of our proposed neural modules, and our reproducible experiment code base and Makefiles. We will also release these materials publicly upon paper acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All relevant hyper-parameters and baselines are disclosed in the main manuscript. The standard splits of the datasets were used.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification:
    • Our results on constructing a Tversky Projection (Section 3.1) that models the xor function is an existence proof, and doesn’t need statistical significance.
    • In our empirical results on learning XOR via gradient descent Appendix D reports the mean and standard error of losses, accuracies and convergence rates across a large sample of 12 9601296012\,96012 960 trained models.
    • Our empirial results with larger neural networks Resnet-50, and GPT2 do not include error bars due to elevated computation costs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Detailed computation resources and times in Appendix H.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed the NeurIPS code of ethics, and have not found that our work violates it in any way.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our work makes progress towards psychologically plausible neural networks that are naturally interpretable by humans, which is a positive societal impact. We do not explicitly discuss potentially negative societal impacts because such discussion would be speculative. However we do recognize that all scientific artifacts can potentially lead to intended or inadvertent societal harms.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: To the best of our knowledge, our research artifacts do not have an immediate high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: This work uses standard, widely used publicly available datasets, which were appropriately cited in the manuscript.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: This paper introduces new neural network building blocks implemented in PyTorch. Our supplementary materials will include code and appropriate documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: No crowd-sourcing experiment or human subjects were used in this research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not850 involve LLMs
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@ 4 Discussion
Our language modeling experiments suggest the existence of an optimal feature bank size, though determining this value a priori remains an open question. In the XOR setting, we observed that Tversky networks can be sensitive to hyperparameters, underscoring the importance of standard tuning practices. For our vision experiments, no tuning was performed; our goal was simply to demonstrate the existence of Tversky models that can match or surpass baseline accuracy. We encourage practitioners to apply standard hyperparameter tuning to further improve performance, especially since Tversky networks are a novel architecture, whereas the baselines we compare against have benefited from years of community-driven refinement.
Weight tying in language models is adequate. Our experiments with GPT-2 on PTB showed that weight tying can be beneficial in both language modeling heads employing Tversky and linear projections. Our interpretation of projection layers as measurement of similarity of stimuli to prototypes explains the adequacy of weight tying in language models: Token embeddings are prototypes, and the language modeling head classifies output tokens by measuring their similarity to those prototypes. This practice with Tversky projection layers is further principled given the underlying theory of similarity.
Data-domain projection parameter specification could enable interpretable editing. For instance, excluding handwritten sevens with the mid horizontal stroke from VisualTverskyMNISTNet’s training set would result in a prototype without that stroke, and high error rates on handwritten sevens with that stroke. The resulting model could be edited by adding a feature specifying the stroke, and editing the prototype to include the stroke. This neural network editing approach could enable the understanding and eradication of dangerous model biases.
@@ Appendix B Illustration of Data-Domain Visualization of Projection Layers
Appendix C Training a Tversky Projection to Learn XOR with Gradient Descent
C.1 Task and Data
@@
C.2 Method
To empirically estimate the sensitivity of Tversky projection layer’s convergence to its hyperparameters, we train 12 96012\,960 xor models consisting of the following combinations of hyperparameters:
@@ C.3 Results
Tversky projection layer successfully modeled xorxor with as little as 1 feature. $p(conv)$ was maximal with 16 features, but not monotonously increase with feature count.
@@
Figure 6: Example of learned TverskyProjection layers optimized to model the xor function. While multiple solutions are possible (top row), some initializations do not lead to convergence (botton row). The decision boundaries were drawn as follows. Random input vectors [x,y] are uniformly sampled in the range [-2,-2] to [2,2] the tip of each vector is ploted as a colored dot, with the color representing the trained Tversky Projection layer’s output. Tversky projection layers modeling the xorxor function, which has a boolean domain and range extend xorxor to the real vector space. Successful models should show data points [0,0] and [1,1] in purple (class 0), and [0,1] and [1,0] in yellow (class 1)
@@
Table 4: Mean $\pm$ std error of loss and accuracy for TverskyXORNet with 6 intersection reduction methods ($A\cap B$) and 2 difference reduction methods (ABA-B). Each row corresponds to the aggregate of 3x3x2x6x9 = 972 training sessions with: 3 fbank and prototype initialization methods(orthogonal, normal, uniform), 2 normalization modes (normalized, and not normalized), 6 feature bank sizes (1, 2, 4, 8, 16, 32), and 9 random seeds. In each case, models were trained for 1000 epochs
$A\cap B$ ABA-B n loss acc best acc p(conv)
@@
gmean ignorematch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $0.00\pm 0.00$
gmean substractmatch 972 $nan\pm nan$ $0.50\pm 0.00$ 0.5 $0.00\pm 0.00$
@@
Appendix D Language Modeling with GPT-2 and Tversky-GPT-2 on PTB
D.1 Initialization: OpenAI Released Weights
@@ D.2 Initialization: Random Weights
Appendix E Examples of Semantic Fields in TverskyGPT2’s Set-Centric Token Representation
Figures  8 ,9 ,10 ,11 and 12 show examples of visualizations of semantic fields formed using set expressions on TverskyGPT2 tokens represented as sets. For each token, the top 1000 features (sorted by dot-product with the token) are considered, and represented as the smallest colored circles. Tokens are represented as yellow circles, and connected to the features they comprise with grey lines. All distinct intersections of features are colored in the same color, and connected to a circle of the same color representing a distinctive semantic field. Semantic fields of interest are annotated with a text box listing the top 50 tokens by semantic score (See Section 3.4) in that semantic field, permitting its visualization. A force-directed graph layout is employed to position the circles corresponding to tokens, features, and semantic fields. Note that the Ġ character (U+0120) represents space in GPT2 tokenization, distinguishing tokens that appear at the beginning of words from other tokens.
Figure 8: Visualization of the semantic fields formed by the distinctive features of related tokens. Notice that TverskyGPT2 tokens capture semantics related to orthographic markings, and morphological function. The kingking token (as in baking) is related to other suffixes such as kedked, kerker and keskes (as in baked, baker and bakes)
Figure 9: Visualization of the semantic fields formed by the distinctive features of tokens likelike (likelovelike-love) and lovelove (lovelikelove-like). likelike, distinctively from lovelove, which also captures the sense of sentiment, captures the sense of similarity, with tokens such as akin, alike, unlike and reminiscent.
Figure 10: Visualization of the semantic fields formed by the distinctive features of tokens gogo (gogonewentgoinggo-gone-went-going), gonegone (gonegowentgoinggone-go-went-going), wentwent (wentgonegogoingwent-gone-go-going), and goinggoing (goinggonewentgogoing-gone-went-go), respectively specifying the concept of verb forms infinitive, past participle, past tense and present participle.
Figure 11: Visualization of the semantic fields formed by the distinctive features of the common features of tokens badbad and goodgood ($bad\cap good-worse-better-worst-best$), worseworse and betterbetter ($worse\cap better-bad-good-worst-best$), and worstworst and bestbest ($worst\cap best-bad-good-worse-better$). These semantic fields algebraically specify the concepts of adjective, comparative, and superlative.
Figure 12: Visualization of the semantic fields capturing the distinctive senses of close synomyms coachcoach, teacherteacher, mentormentor, and guideguide. Notice that the distinctive features of coachcoach (coachteachermentorguidecoach-teacher-mentor-guide) shows terms related to sport such as trainer, teammate, Football and ESPN whereas the distinctive features of teacherteacher (teachercoachmentorguideteacher-coach-mentor-guide) show terms related to formal education such as educator, classroom, instructor, lecturer and professor.
Figure 13: Visualization of the semantic fields capturing the distinctive feature of tokens manman, womanwoman, kingking and queenqueen, and the distinctive semantics of the common features of manman and kingking, manman and womanwoman, womanwoman and queenqueen, queenqueen and kingking, and manman and womanwoman
Appendix F [Visual]MNISTNet, and [Visual]TverskyMNISTNet,
Two neural networks, VisualMNISTNet, and VisualTverskyMNISTNet (Figure 14) are trained to perform MNIST handwritten digit classification. The two neural network architectures share the same convolutional feature extraction stack yielding a 36-dimensional vector representation of the input image. They differ in how those vectors are projected onto the 10-dimensional output vectors corresponding to 10 digit classes. VisualMNISTNet employs a stack of 120, 84, and 10 unit linear projection layers following LeCun et al. [1998]. However, the ReLU activation function is used instead of the logistic sigmoid. VisualTverskyMNISTNet employs a single Tversky Projection layer with 10 prototypes and 20 features. Both neural networks employ the method described in Section 2.5, facillitating our qualitative analysis of the learned prototypes and features.
Model Params Valid ACC
MNISTNet 21 394 99.1%
VisualMNISTNet 28 394 99.1%
TverskyMNISTNet 7 023 98.7%
VisualTverskyMNISTNet 29 463 98.7%
Figure 15: Learned visual prototypes in the baseline and tversky variants of the neural network described in Figure 14 after epochs 0, 200, 400 and 600.
Figure 16: Two MNIST digits: a two (A), and a one (B). A grid of top-9 instances ranked by semantic score in the ABA-B, $A\cap B$ and BAB-A semantic fields are shown (see Section 3.4). The distinctive features of A exhibit curviness. The distinctive features of B exhibit a vertical stroke slanted to the right.
@@ H.1 Resnet-50 / NABirds Experiments
H.2 GPT2 Experiments